OpenStudy (aravindg):

If f(x) = loge x and x >0, prove that f( u v w) = f(u) + f(v) + f(w).

OpenStudy (anonymous):

log base (e)?

OpenStudy (anonymous):

so ln(x)

OpenStudy (aravindg):

ya log base e

OpenStudy (anonymous):

f(u*v*w)=ln(u*v*w)=ln(u)+ln(v)+ln(w)

OpenStudy (anonymous):

rules of log

OpenStudy (aravindg):

what is ln??

OpenStudy (anonymous):

f[u*v*w]=log(u*v*w)=log(u)+log(v)+log(w) = f[u]+f[v]+f[w]

OpenStudy (zarkon):

\[\ln(x)=\log_e(x)\]

OpenStudy (anonymous):

and f(u)=ln(u) f(v)=ln(v) f(w)=ln(w) f(u)+f(v)+f(w)=ln(u)+ln(v)+ln(w)

OpenStudy (anonymous):

and as zarkon wrote ln means natural log or a log with the base value being e

OpenStudy (aravindg):

is log a +log b = log a *log b???

OpenStudy (anonymous):

log(a)+log(b)=log(a*b)